# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v4

@@ 1.4 Limitations
Like any empirical analysis of its kind, also our study makes assumptions and faces some limitations. First, a central assumption that all such studies need to make for a split of the price zone are the cross-zonal line capacities. Transmission capacity allocation in the field is a complex task with several decision made by TSOs, that cannot easily be predicted and modeled (see Section 2.2.1). In line with earlier studies (see Section 2.2.1 for a discussion), we compute market clearing solutions in which only 80% of the transmission line limits are considered to take into account security considerations. With higher security margins there might be higher differences in the prices between zones. Second, as in any simulation, the way how we compute redispatch costs can only be an approximation of real-world practices, which are complex (see decision BK8-18-0007-A by the German Bundesnetzagentur). For example, we do not provide a detailed accounting of curtailment costs. We also assume that every generator can be redispatched, but this is not the case in reality. So, the computed redispatch costs can only provide a lower bound, but we argue that it provides a fair comparison. Third, our examination focuses on the German day-ahead market without incorporating neighboring countries, cross-border trades, or loop flows through other countries. We do incorporate 100 neighbouring nodes not in Germany though to mitigate the effect, and argue that the consideration of more nodes outside Germany will not lead to higher differences in the prices between zones in Germany but rather the opposite. Fourth, we do not consider investment incentives and the long-run effects of price differences. Such questions are considered in multi-level game-theoretical models, but these incentives also contribute to more similar prices rather than price differences across the zones (Ambrosius et al.,, 2020). Fifth, we assume fixed demand from the ENTSO-E dataset, neglecting the potential for demand response under nodal prices. Incentives for demand response would be in favour of nodal pricing rules. Sixth, the current European market is characterized by portfolio bidding, differing from the unit-commitment bids underlying both the study by ENTSO-E and our analysis. Portfolio bids allow optimization of trades across assets and differ from those feasible in nodal market designs. Incorporating such considerations into our analysis would require making strong and potentially unwarranted assumptions regarding market participant behavior. Neither the LMP study nor we make such assumptions. Seventh, as any other study of this sort, we do not model the intraday or forward markets. Finally, the data set provided by ENTSO-E was constructed for the target year 2025, and it is important to acknowledge that analyses can vary based on differing assumptions regarding supply and demand. However, the ENTSO-E dataset represents the best source available today for such a study. The assumptions made for our analysis are described in Section 3.
@@
We also want to emphasize that the discussion surrounding zonal and nodal pricing extends beyond dispatch and prices. In the ongoing Bidding Zone Review, a comprehensive approach is adopted which considers factors such as liquidity, security of supply, transition costs, and investments in low-carbon technologies, which we can not replicate in our analysis.
@@ 1.5 Organization
The remainder of this article is structured as follows. In Section 2, we introduce our unit commitment model and formalize nodal and zonal market clearing. We also review the IP, CH, and Join pricing rules. Section 3 elaborates on the data sources used to implement our dispatch model and outlines our experimental design. Section 4 presents the key results of our analysis of different market clearing and pricing rules. Lastly, Section 5 provides a summary and conclusions.
2 Clearing and Pricing on Electricity Markets
2.1 Preliminaries
We consider an electricity market consisting of a set of buyers $B$ and a set of sellers $S$, each located at nodes $N$ in an interconnected electricity network. The set of power lines $L$ is encoded as pairs of nodes $(n,m)$, and we consider multiple periods $T$. An item in this market corresponds to a unit of electricity at a particular node $n\in N$ at a specific time $t\in T$.
A buyer $b\in B$ possesses a valuation function $v_{b}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}$ that quantifies the buyer’s preferences and constraints. Similarly, a seller $s\in S$ has a cost function $c_{s}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}$.
The market operator collects these buy and sell bids to calculate a feasible allocation. As electricity is transmitted over the network, the allocation is subject to physical power flows, encoded as a constraint set $\Psi$. An accurate representation of the transmission network would require $\Psi$ to be equivalent to AC optimal power flow (ACOPF) constraints (Molzahn & Hiskens,, 2019). However, it is well known that the ACOPF is intractable for realistic problem sizes, and electricity market operators need to employ computationally scalable approximations of the ACOPF.
2.2 Market Clearing
In practice, market operators employ different bidding languages to encode $v_{b}$ and $c_{s}$, as well as different approximations $\Psi$ of the power flow constraints. In this work, we use the data released for the European bidding zone review to compare allocation and prices under different market clearing mechanisms.
In full generality, the market operator seeks an allocation $(x,y)$, where $x=(x_{b})_{b\in B}$, $x_{b}\in\mathbb{R}^{N\times T}$ is the allocation vector of buyers and $y=(y_{s})_{s\in S}$,$y_{s}\in\mathbb{R}^{N\times T}$ is the allocation vector of sellers. The market operator identifies an allocation that maximizes welfare, by solving the following optimization model:
@@ 2.2 Market Clearing
On European day-ahead markets, buyers and sellers submit hourly bids and block orders (NEMO Committee,, 2019). The BZR data, however, does not contain bids of this structure. Instead, it provides generator and load characteristics similar to unit commitment problems in U.S. markets. Specifically, for a buyer $b$ located in a node $n_{b}$ [^7] For $t\in T,n\in N\setminus\{n_{b}\},x_{b,nt}=0$. there is only information on a fixed demand profile $P_{b}\in\mathbb{R}^{N\times T}$, while price-elastic bids are unavailable. Consequently, the valuation function of $b$ simplifies to strict demand satisfaction.
$v_{b}(x_{b})=\begin{cases}-\infty,\;\;x_{b}\neq P_{b}\\ 0,\;\;\;x_{b}=P_{b}\end{cases}$ (2)
@@
For sellers/generators $s\in S$, the BZR data release includes more detailed information. Each generator $s$ has a minimum and maximum production quantity $\underline{P}_{s}\in\mathbb{R}$ and $\overline{P}_{s}\in\mathbb{R}$ at their node $n_{s}$. Moreover, once a generator has been started, there is a minimum uptime constraint to run for at least $\underline{R}_{s}\in\mathbb{Z}_{0}^{+}$ periods. Generators incur variable costs $g_{s}$ to produce electricity and fixed costs $h_{s}$ whenever the generator runs. We use a binary commitment variable $u_{st}\in\{0,1\}$ to model a generator’s constraint set $P_{s}$.
$P_{s}=\{y_{s}\in\mathbb{R}^{N\times T}:\;$ $y_{s,n_{s}t}\geq\underline{P}_{st}u_{st}$ $\;\forall t\in T,$
$y_{s,n_{s}t}\leq\overline{P}_{st}u_{st}$ $\;\forall t\in T,$
$\phi_{s,n_{s}t}\geq u_{st}-u_{s(t-1)}$ $\;\forall t\in T\setminus\{T_{0}\},$
$\sum_{i=t-\underline{R}_{s}+1}^{t}\phi_{s,n_{s}i}\leq u_{st}$ $\;\forall t\in T\setminus\{T_{0}\},$
$y_{s,nt}=0$ $\;\forall n\in N\setminus\{n_{s}\},t\in T$
$u_{st}\in\{0,1\}$ $\;\forall s\in S,t\in T.\}$
With this constraint set, we define a generator’s cost function $c_{s}$ as follows.
$c_{s}(y_{s})=\begin{cases}\infty,\;\;y_{s}\notin P_{s}\\ \sum_{t\in T}g_{s}y_{s,n_{s}t}+\sum_{t\in T}h_{s}u_{st},\;\;y_{s}\in P_{s}\end{cases}$ (3)
Given that only price-inelastic demand bids are available, the objective of problem (1) can be interpreted as minimizing generators’ costs required to fulfill the demand.
@@
During the 2000s and 2010s, many U.S. ISO markets moved from zonal to nodal pricing. Recognizing that an accurate representation of the transmission network in the form of the ACOPF is computationally infeasible, market operators resort to linearized versions of power flow. The most common implementation is the Direct Current OPF (DCOPF), which is based on three simplifying assumptions (Stott et al.,, 2009; Molzahn & Hiskens,, 2019), i.e., (1) uniform voltage magnitudes, (2) negligibly small voltage angle differences, and (3) neglecting line resistances and reactive power. Under these assumptions, the flow constraints can be expressed as simple linear constraints.
$\Psi^{DC}=\{(x,y):\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall n\in N,t\in T\}$ (4)
@@
In (4), $B_{nm}$ is the susceptance of the line connecting nodes $n$ and $m$. The voltage angle decision variable for node $n$ and period $t$ is denoted as $\theta_{nt}$. If such a line does not exist, then $B_{nm}=0$, and $B_{nn}=0\;\forall n$. The constraint set in (4) can be enriched by further constraints, such as box constraints on the voltage angles, and more (Li et al.,, 2022). We use such a standard DCOPF model in our analysis, similar to what was used in the ENTSO-E study. Note that this representation of the DCOPF using line susceptances is equivalent to the representation that considers PTDFs. When the representation $\Psi$ of the transmission network includes all nodes in the transmission network, we refer to nodal market clearing. When only a subset of the grid is considered, we speak of zonal market clearing, as discussed in the following. Power system operators need to consider contingencies for a secure operation of the network. The system should satisfy the ($N-1$) criterion to guarantee that the power system is operated safely in the normal condition or in any contigency case (Christie et al.,, 2000; Marien et al.,, 2013; Weinhold & Mieth,, 2024). Rather than a full security-constrained optimal power flow (SCOPF) model (Hinojosa,, 2020), we consider a safety margin of 20% on all transmission lines.
2.2.1 Zonal Market Clearing
@@
The key idea of zonal pricing is to aggregate nodes into zones and only consider transmission flows across zones. This simplifies the constraint set $\Psi$ significantly, enhances computational scalability, and allows for portfolio bidding of larger market participants. An important parameter of a zonal power system is the determination of the capacities of interconnectors between the zones.
@@
In the European market, two approaches have been implemented: the net transfer capacity (NTC) model and the flow-based market coupling (FBMC) methodology (NEMO Committee,, 2019). FBMC was introduced in Central Western Europe in 2015 and has since been extended to neighboring markets. In FMBC, the net export possible from one zone to another is limited by the remaining available margin (RAM) on the critical network elements connecting the zones. The RAM is determined by the maximum allowable flow (e.g., thermal capacity) subtracting the physical flow before the optimization (the base case), and a flow reliability margin to cover uncertainty in the capacity and allocation computations (Marien et al.,, 2013; Weinhold & Mieth,, 2024). The Flow-Based Market Coupling with Generation Shift Keys (FBMC-GSK) is the specific methodology currently being used in the EU. The way how the interconnection capacity between two price zones is calculated and allocated in FBMC-GSK depends on how various parameters are set and the critical network elements considered by the TSOs (Aravena et al.,, 2021), and cannot be predicted precisely for an eventual split of the German price zone. Furthermore, the FBMC approach necessitates access to a PTDF matrix that indicates how zone net position changes affect cross-zonal flows. Such a PTDF matrix is not included in the BZR data.
@@
In our zonal market clearing implementation, we consider all interconnectors as critical network elements and model their physical characteristics via the DCOPF constraints (4). In line with earlier literature (Tiedemann et al.,, 2024; Wyrwoll et al.,, 2018; Voswinkel et al.,, 2019), we compute a base case that only takes 80% of the capacities of all interconnectors into account.Similar to the nodal model, this should take into account the $(N-1)$ security criterion (Christie et al.,, 2000). The set of feasible zonal solutions can be written as:
$\Psi^{Zonal}=\{(x,y):\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{n\in Z,m\notin Z}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall z\in Z,t\in T\}$ (5)
The mapping of nodes to zones is suggested by the ENTSO-E study on locational marginal pricing (ACER, 2022b, ). Four different zonal splits were proposed for the German bidding zone, suggesting between two to four price zones.
2.2.2 Redispatch
@@
Current redispatch mechanisms consider two paradigms: Firstly, cost-based redispatch, as implemented in Germany, reimburses redispatched generators for their additional costs or lost profits and makes them indifferent to the previous market outcome. Secondly, market-based redispatch aims at a bid-based procurement of redispatch volumes. Market-based redispatch was criticized as prone to market power abuse and causing inc-dec gaming, i.e., strategic overbidding in day-ahead markets to generate higher profits in subsequent redispatch markets (Hirth & Schlecht,, 2020).
We simulate a cost-based redispatch. As redispatch should be cost-minimal, the approach is to find a modified allocation that satisfies the constraint set of $\Psi^{DC}$ and, at the same time, minimizes the required compensations to sellers for their changes in generation. Demand is excluded from redispatch, as it is current practice in Germany. More formally, with $(x^{Zonal},y^{Zonal})$ as the outcome of the zonal allocation and $f_{nmt}=B_{nm}(\theta_{nt}-\theta_{mt})$ obtained from (5) such that nodes $n$ and $m$ belong to two different zones, we solve the following problem:
$\min_{x,y}\quad$ $\sum_{s\in S}\lvert c_{s}(y_{s})-c_{s}(y_{s}^{Zonal})\rvert$ (6)
subject to $B_{nm}(\theta_{nt}-\theta_{mt})\leq f_{nmt}\ \ \forall n,m\in N,zone(n)\neq zone(m),t\in T$
$(x,y)\in\Psi^{DC}$
Here, $zone(n)$ denotes the zone to which node $n$ belongs. The objective of the optimization problem described in (6) is to adjust the zonal allocation $(x^{Zonal},y^{Zonal})$ in a manner that yields a DCOPF-feasible solution $(x,y)$ at minimal costs. This adjustment involves compensating sellers for any changes in their dispatched quantities during redispatch. For each seller $s$, this compensation is determined by the difference between the costs they incur under the initial zonal allocation $c_{s}(y_{s}^{Zonal})$ and the costs under the final DCOPF-feasible allocation $c_{s}(y_{s})$. This assumes that every generator can be redispatched, which might not be the case in practice. Note that the constraints on cross-zonal line capacities from the zonal model are also taken into account in the redispatch computation. The objective of (6) is to minimize the total costs associated with these adjustments. As part of our experiments we also report results for redispatch that minimizes volumes instead of costs, i.e., the redispatch objective function changes to $\min_{x,y}\;\sum_{s\in S}\lvert y_{s}-y_{s}^{Zonal}\rvert$.
@@
Problem (6) yields the final DCOPF-feasible allocation $(x,y)$ that can be implemented as physical power flows. Note that a DCOPF-feasible solution may still be ACOPF-infeasible and require further adjustments to the allocation, which we disregard in this work. Market participants subject to redispatch must be compensated for lost profits, reducing overall welfare. In our numerical experiments, we thus consider system costs as the total cost of generation before redispatch, redispatch costs themselves, and total costs as the sum of system and redispatch costs.
2.3 Pricing on Non-Convex Markets

### v6

@@ 1.6 Limitations
Like any model-based analysis, our study makes assumptions and faces some limitations. First, a central assumption that all such studies need to make for a BZ split are the cross-zonal line capacities. Transmission capacity allocation is a complex task with several decision made by TSOs, which cannot easily be predicted and modeled (see Section 2.1.2). In line with earlier studies (see Section 2.1.2), we compute market clearing solutions in which only 80% of the transmission line limits are considered to take into account security considerations. With higher security margins there might be higher differences in the prices between zones, and also lower redispatch costs. Second, as in any simulation, the way how we compute redispatch can only be an approximation of real-world practices, which are complex (see decision BK8-18-0007-A by the German Bundesnetzagentur). For example, we assume that every generator can be redispatched, but this is not the case in reality. Moreover, we do not consider countertrade measures and restrict our analysis to day-ahead redispatch. So, the computed redispatch costs serve only as an approximation of the redispatch costs encountered by TSOs, but we argue that they provide a fair comparison. Third, our examination focuses on the German day-ahead market without incorporating neighboring countries, cross-border trades, or loop flows through other countries. We do incorporate 100 neighboring nodes not in Germany though to mitigate the effect. However, flows from neighboring countries could enhance congestion and, thus, lead to higher redispatch volumes and costs in Germany. Fourth, we do not consider the long-run effects of the prices and congestion management regime corresponding to specific market clearing models. While this paper focuses on short-run inefficiencies, long-run inefficiencies – such as suboptimal location choices and technology decisions for generation and consumption – are also crucial factors to consider when evaluating a market clearing model. Fifth, we assume fixed, inelastic demand from the ENTSO-E dataset [^10] According to NEMO Committee, (2024), demand is inelastic and must be satisfied in any feasible solution., neglecting the potential for demand response under nodal prices. Incentives for demand response would be in favor of nodal pricing rules. Moreover, we assume that the demand cannot be redispatched, as it is current practice in Germany. Sixth, the current European market is characterized by portfolio bidding, differing from the unit-commitment bids underlying both the study by ENTSO-E and our analysis. Portfolio bids allow optimization of trades across assets and differ from those feasible in nodal market designs. Incorporating such considerations into our analysis would require making strong and potentially unwarranted assumptions regarding market participant behavior. Neither the LMP study nor we make such assumptions. Seventh, as any other study of this sort, we do not model the intraday or forward markets. Finally, the data set provided by ENTSO-E was constructed for the target year 2025 based on input data and assumptions collected in 2019, and might not reflect the state of the electricity market in future years. It is also important to acknowledge that analyses can vary based on differing assumptions regarding supply, demand and electricity grid, as well as the employed methods. The data assumptions made for our analysis are described in B. We also highlight the modelling differences compared to the BZR study in E.
@@
We also want to emphasize that the discussion surrounding zonal and nodal pricing extends beyond dispatch and prices. In the BZR, a comprehensive approach is adopted which considers factors such as liquidity, security of supply, transition costs, and investments in low-carbon technologies, which we can not replicate in our analysis.
@@ 1.7 Organization
The remainder of this article is structured as follows. In Section 2.1, we introduce our unit commitment model and formalize nodal and zonal market clearing. We also review the IP, CH, and Join pricing rules. Section 2.2 outlines our experimental design. Section 3 presents the key results of our analysis. Lastly, Section 4 provides a summary and conclusions.
2 Methods
2.1 Clearing and Pricing on Electricity Markets
2.1.1 Preliminaries
We consider an electricity market consisting of a set of buyers $B$ and a set of sellers $S$, each located at nodes $N$ in an interconnected electricity network. The set of power lines $L$ is encoded as pairs of nodes $(n,m)$, and we consider multiple periods $T$. An item in this market corresponds to a unit of electricity at a particular node $n\in N$ at a specific time $t\in T$. A buyer $b\in B$ possesses a valuation function $v_{b}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}$ that quantifies their preferences. Similarly, a seller $s\in S$ has a cost function $c_{s}:\mathbb{R}^{N\times T}\rightarrow\mathbb{R}_{\geq 0}$.
The market operator collects these buy and sell bids to calculate a feasible allocation. As electricity is transmitted over the network, the allocation is subject to physical power flows, encoded as a constraint set $\Psi$. An accurate representation of the transmission network would require $\Psi$ to be equivalent to AC optimal power flow (ACOPF) constraints (Molzahn & Hiskens, , 2019). However, it is well known that the ACOPF is intractable for realistic problem sizes, and market operators need to employ computationally scalable approximations of the ACOPF.
2.1.2 Market Clearing
In practice, market operators employ different bidding languages to encode $v_{b}$ and $c_{s}$, as well as different approximations $\Psi$ of the power flow constraints. In full generality, the market operator seeks an allocation $(x,y)$, where $x=(x_{b})_{b\in B}$, $x_{b}\in\mathbb{R}^{N\times T}$ is the allocation vector of buyers and $y=(y_{s})_{s\in S}$, $y_{s}\in\mathbb{R}^{N\times T}$ is the allocation vector of sellers. The market operator identifies an allocation that maximizes welfare by solving the following optimization model:
@@ 2.1.2 Market Clearing
On European day-ahead markets, buyers and sellers submit hourly bids and block orders (NEMO Committee, , 2019). The BZR data, however, does not contain bids of this structure. Instead, it provides generator and load characteristics similar to unit commitment problems in U.S. markets. Specifically, for a buyer $b$ located in a node $n_{b}$ there is only information on a fixed demand profile $P_{b}\in\mathbb{R}^{N\times T}$, while price-elastic bids are unavailable. Thus, the valuation function of $b$ simplifies to strict demand satisfaction, i.e., $v_{b}(x_{b})=0$ if $x_{b}=P_{b}$ and $v_{b}=-\infty$ otherwise.
@@
For sellers/generators $s\in S$, the BZR data release includes more detailed information. Each generator $s$ has a minimum / maximum production quantity $\underline{P}_{s}$ / $\overline{P}_{s}\in\mathbb{R}_{\geq 0}$ at their node $n_{s}$. Moreover, once a generator has been started, it must run for at least $\underline{R}_{s}\in\mathbb{Z}_{0}^{+}$ periods. A generator incurs variable costs $g_{s}$ and fixed costs $h_{s}$, and its cost function is $c_{s}(y_{s})=\sum_{t\in T}g_{s}y_{s,n_{s}t}+\sum_{t\in T}h_{s}u_{st}$. Given that only price-inelastic demand bids are available, the objective of problem (1) can be interpreted as minimizing generators’ costs required to fulfill the demand.
@@
During the 2000s and 2010s, many U.S. ISO markets moved from zonal to nodal pricing. Recognizing that an accurate representation of the transmission network in the form of the ACOPF is computationally infeasible, market operators resort to linearized versions of power flow. The most common implementation is the Direct Current OPF (DCOPF), which is based on three simplifying assumptions (Stott et al., , 2009; Molzahn & Hiskens, , 2019), i.e., (1) uniform voltage magnitudes, (2) negligibly small voltage angle differences, and (3) neglecting line resistances and reactive power. Under these assumptions, the constraint set $\Psi$ can be expressed as simple linear constraints:
$\Psi^{DC}=\{(x,y,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\theta}):\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall n\in N,t\in T\}.$ (2)
The following MILP encodes the nodal clearing problem:
$\min\limits_{x,y,u,\theta}\quad$ $\sum_{s\in S}c_{s}(y_{s})$ (Nodal Clearing)
subject to $\sum_{s\in S}y_{s,nt}-\sum_{b\in B}x_{b,nt}=\sum_{m\in N}-B_{nm}(\theta_{nt}-\theta_{mt})$ $\forall n\in N,t\in T$ (3)
$\underline{F}_{nm}\leq B_{nm}(\theta_{nt}-\theta_{mt})\leq\overline{F}_{nm}$ $\forall n,m\in N,t\in T$ (4)
$\underline{P}_{st}u_{st}\leq y_{s,n_{s}t}\leq\overline{P}_{st}u_{st}$ $\;\forall s\in S,t\in T$ (5)
$y_{s,nt}=0$ $\;\forall s\in S,n\in N\setminus\{n_{s}\},t\in T$ (6)
$\phi_{st}\geq u_{st}-u_{s(t-1)}$ $\;\forall s\in S,t\in T\setminus\{T_{0}\}$ (7)
$\sum_{i=t-\underline{R}_{s}+1}^{t}\phi_{si}\leq u_{st}$ $\;\forall s\in S,t\in T\setminus\{T_{0}\}$ (8)
$c_{s}(y_{s})=\sum_{t\in T}g_{s}y_{s,n_{s}t}+\sum_{t\in T}h_{s}u_{st}$ $\forall s\in S$ (9)
$x_{b,n_{b}t}=P_{bt}$ $\forall b\in B,t\in T$ (10)
$x_{b,nt}=0$ $\forall b\in B,n\in N\setminus\{n_{b}\},t\in T$ (11)
$\theta_{n^{*}t}=0$ $\forall t\in T$ (12)
@@
A summarizes the notation. We use a standard DCOPF model in our analysis, similar to what was used in the ENTSO-E study. Note that power system operators need to consider contingencies for a secure operation of the network. In particular, the system should satisfy the ($N-1$) criterion to guarantee that the power system is operated safely in the normal condition or in any contigency case (Christie et al., , 2000; Marien et al., , 2013; Weinhold & Mieth, , 2024). Rather than a full security-constrained optimal power flow (SCOPF) model (Hinojosa, , 2020), we consider a safety margin of 20% on all transmission lines.
Zonal Market Clearing
@@
The key idea of zonal pricing is to aggregate nodes into zones and only consider transmission flows across zones, which simplifies the constraint set $\Psi$ and enhances computational scalability. An important parameter of a zonal system is represented by the capacities of the cross-zonal lines.
@@
In the European market, two approaches have been implemented: the net transfer capacity (NTC) model and the flow-based market coupling (FBMC) methodology (NEMO Committee, , 2019). FBMC was introduced in Central Western Europe in 2015 and has since been extended to neighboring countries. In FMBC, the net export possible from one zone to another is limited by the remaining available margin (RAM) on the critical network elements (CNEs) connecting the zones. The RAM is determined by the maximum allowable flow (e.g., thermal capacity) subtracting the physical flow before the optimization (the base case), and a flow reliability margin to cover uncertainty in the capacity and allocation computations (Marien et al., , 2013; Weinhold & Mieth, , 2024). FBMC with Generation Shift Keys (FBMC-GSK) is the specific methodology currently being used in the EU. The way how cross-zonal capacity is calculated and allocated in FBMC-GSK depends on how various parameters are set and the CNEs considered by the TSOs (Aravena et al., , 2021), and cannot be predicted precisely for an eventual split of the German BZ.
@@
In our zonal market clearing implementation, we consider all interconnectors as critical network elements and model their physical characteristics via the DCOPF constraints (2). In line with earlier literature (Tiedemann et al., , 2024; Wyrwoll et al., , 2018; Voswinkel et al., , 2019), we compute a base case that only takes 80% of the capacities of all interconnectors into account. Similar to the nodal model, this should take into account the $(N-1)$ security criterion (Christie et al., , 2000). The zonal flow constraints can be written as:
$\Psi^{Zonal}=\{(x,y,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\theta}):\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{n\in z,m\notin z}-B_{nm}(\theta_{nt}-\theta_{mt})\;\forall z\in Z,t\in T\}$ (13)
The zonal clearing problem is formulated as follows:
$\min\limits_{x,y,u,\theta}\quad$ $\sum_{s\in S}c_{s}(y_{s})$ (Zonal Clearing)
subject to $\sum_{s\in S}y_{s,zt}-\sum_{b\in B}x_{b,zt}=\sum_{n\in z,m\notin z}-B_{nm}(\theta_{nt}-\theta_{mt})$ $\forall z\in Z,t\in T$ (14)
$\underline{F}_{nm}\leq B_{nm}(\theta_{nt}-\theta_{mt})\leq\overline{F}_{nm}$ $\forall n,m\in N,z(n)\neq z(m),t\in T$ (15)
$\theta_{z^{*}t}=0$ $\forall t\in T$ (16)
[…]
We denote by $(x^{Zonal},y^{Zonal},u^{Zonal},\theta^{Zonal})$ the optimal solution of the zonal clearing problem.
Redispatch
@@
In the Single Day-Ahead Coupling Region (SDAC), the economic allocation is determined using the Euphemia algorithm (NEMO Committee, , 2024). This allocation includes decisions on order acceptance and power flows between the bidding zones. The allocation results are published around 12:30 CET [^11] The order book closes at 12:00 CET and then Euphemia computes a solution within 12 minutes., setting fixed day-ahead schedules. After these results are released, redispatch begins, which occurs during the period leading up to real-time delivery.
@@
Current redispatch mechanisms consider two paradigms: Firstly, cost-based redispatch, as implemented in Germany, reimburses redispatched generators for their additional costs or lost profits and makes them indifferent to the previous market outcome. Secondly, market-based redispatch aims at a bid-based procurement of redispatch volumes. Market-based redispatch was criticized as prone to market power abuse and causing inc-dec gaming, i.e., strategic bidding in day-ahead markets to generate higher profits in subsequent redispatch markets (Hirth & Schlecht, , 2020).
The precise redispatch methodology implemented by TSOs is not publicly disclosed. While Article 35 of the CACM Regulation states that “Each TSO shall coordinate the use of redispatching and countertrading resources, taking into account their impact on operational security and economic efficiency”, it does not detail the coordination process or provide formal definitions of operational security and economic efficiency. The German legislation (§13 EnWG) states that “In the case of power and voltage-related adjustments to active power generation or active power consumption […], among several suitable measures […], those measures are to be selected which are expected to cause the lowest overall costs”. However, the legislation does not provide the mathematically precise definition that has to be considered by TSOs. There are several studies that touch upon redispatch computation and employ different models. Ambrosius et al., (2022); Grimm et al., (2019) and Trepper et al., 2015b consider a redispatch problem that has as objective minimizing the cost difference between the redispatch and the spot (zonal) solutions. Moreover, Kunz & Zerrahn, 2015a implement a model that minimizes the sum of redispatch costs, costs associated with load shedding and costs associated with renewable generation curtailment.
In this paper, we consider a minimum-cost redispatch method in line with earlier literature. We assume that demand cannot be redispatched, as it is current practice in Germany. This means that the (inelastic) demand $P_{b}$ for all $b\in B$ will be fully satisfied both in the zonal and in the redispatch solution. Let $R$ denote the set of renewables, $Cur^{res}$ the quantity of renewable generation curtailment corresponding to $res\in R$ and $c_{cur}$ the curtailment cost per MW. Given the outcome of the zonal allocation $y^{Zonal}$, the redispatch problem is:
$\min_{x,y,\theta}\quad$ $\sum_{s\in S}c_{s}(y_{s}-y_{s}^{Zonal})+\sum_{res\in R}c_{cur}Cur^{res}$ (RD-Min-Cost)
subject to $Cur^{res}=\min(0,y^{Zonal}_{res}-y_{res})$ $\forall res\in R$ (17)
[…]
Negative (positive) redispatch occurs when $y_{s}<y^{Zonal}_{s}(y_{s}>y^{Zonal}_{s})$. Negative redispatch implies cost savings since generation costs are avoided, which leads to a decrease in the overall redispatch costs. These costs must be reimbursed by power plants to TSOs. Positive redispatch requires additional generation costs for upward production, causing an increase in the total redispatch costs. This method ensures that the cheap (expensive) generators are selected for upward (downward) redispatch. Note that it also assumes that every generator can be redispatched, which might not be the case in practice. [^12] Until 2021, only production units with a capacity of at least 10 MW could be redispatched. The Grid Expansion Acceleration Act (Netzausbaubeschleunigungsgesetz, NABEG 2.0) introduced the current system, Redispatch 2.0, in which units with a capacity of at least 100 kW can be redispatched. Redispatch 3.0. currently under discussion, aims to further extend it by integrating units with capacities below 100 kW. This would allow for redispatch of all production units, aligning with the method that we implement in this study. We assume that the set of renewables $R$ includes offshore wind, onshore wind and photovoltaic units. The curtailment costs are set to the variable costs presented in Table 15, which reflect the variable and annual operating costs. [^13] Since the curtailment costs are meant to cover the lost revenue plus additional expenses incurred by curtailed renewable units, the costs from Table 15 act as a lower bound for the actual curtailment costs. Note that such costs are not incurred in a nodal market clearing solution (Nodal Clearing problem).
@@
The objective value of problem RD-Min-Cost represents the total redispatch costs. A redispatch solution is interpreted as the final DCOPF-feasible allocation $(x,y,\theta)$ that can be implemented as physical power flows. Note that a DCOPF-feasible solution may still be ACOPF-infeasible and require further adjustments to the allocation, which we disregard in this work. Market participants subject to redispatch must be compensated for lost profits, reducing overall welfare. In our numerical experiments, we thus consider system costs as the total cost of generation before redispatch, redispatch costs themselves, and total costs as the sum of system and redispatch costs.
2.1.3 Pricing on Non-Convex Markets